OpenStudy (anonymous):

What is a polygon called if the sum of its interior angles equals 1440? a dod. dec. or oct?

OpenStudy (mertsj):

Divide 1440 by 180 to see how many triangles the polygon has.

OpenStudy (mertsj):

How many did you get?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

180(n-2) = 1440 180n - 360 = 1440 180n = 1800 n = 10 Therefore, it's dec

OpenStudy (mertsj):

So the number of sides has to be 2 more than the number of triangles. So how many sides does it have?

OpenStudy (anonymous):

petew said it alrdy

OpenStudy (anonymous):

Recall the formula: Total angle in polygon = 180(n-2)

OpenStudy (mertsj):

I saw that. He ruined it for you. Now we know that he can do the problem, but we are not so sure about you.

OpenStudy (anonymous):

I've given the steps to do the problem, all you need to do is understand them.

OpenStudy (anonymous):

no its fine i just really need to remember the definitions to this shapes the formula i wrote it down thx